OpenStudy (anonymous):

If the width of a rectangular piece of carpet is 2x + 1 yards and the length is three yards less than twice as long as the width, what is the area of the carpet? A. 6x² + 2x – 1 B. 8x² + 2x – 1 C. 8x² – 2x – 1 D. 8x² + 6x – 1

OpenStudy (campbell_st):

so the length is \[2 \times(2x + 1) - 3 = 4x + 2 - 3 = 4x -1\] so you are finding the area of a piece of carpet A = (4x -1)(2x + 1) you just need to distribute